Find the change in the volume of 1.0 litre kerosene when it is subjected to an extra pressure of 2.0 × 105 N m-2 from the following data. Density of kerosene = 800 kg m-3 and speed of sound in kerosene = 1330 m s-1.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. We know that v = √K / p Where K = bulk modulus of elasticity ⇒ K = v^2 p = (1330)2 × 800 N/m^2 We know K = (F/A/∆V/V) ⇒ ∆V = Pressure/K = 2 * 10^5/1330 * 1330 * 800 So, ∆V = 0.15 cm^5